You are the operations manager for an airline and you are considering a higher fare level for passengers in aisle seats. How many randomly selected air passengers must you​ survey? Assume that you want to be 99​% confident that the sample percentage is within 3.5 percentage points of the true population percentage. Complete parts​ (a) and​ (b) below.

a. Assume that nothing is known about the percentage of passengers who prefer aisle seats.

n=

​(Round up to the nearest​ integer.)

b. Assume that a prior survey suggests that about 38​% of air passengers prefer an aisle seat.

n=

​(Round up to the nearest​ integer.)
Answer #1

Solution :

Given that,

margin of error = E = 0.035

Z\alpha/2 = 2.576

(a)

\hat p = 0.5

1 - \hat p = 0.5

sample size = n = (Z\alpha / 2 / E)2 * \hat p * (1 - \hat p )

= (2.576 / 0.035)2 * 0.5 * 0.5

= 1355

sample size = n = 1355

(b)

\hat p = 0.38

1 - \hat p = 0.62

= (2.576 / 0.035)2 * 0.38 * 0.62

= 1277

sample size = n = 1277